OpenStudy (dtan5457):

Express 5/3-radical 2 with a rational denominator in the simplest radical form

OpenStudy (dtan5457):

@satellite73

OpenStudy (dtan5457):

@sammixboo

OpenStudy (dtan5457):

@Directrix

OpenStudy (dtan5457):

5 divided by 3-radical 2

OpenStudy (dtan5457):

5 is the numerator, 3- radical 2 on denominator

OpenStudy (jhannybean):

\[\frac{5}{3-\sqrt{2}}\]

OpenStudy (jhannybean):

Since there is a radical in the denominator, you need to take the conjugate of the function on the bottom, that is, if you have \(a+b\) then it's conjugate would be \(a-b\)

OpenStudy (dtan5457):

so i multiply by conjugate

OpenStudy (dtan5457):

would that be 3+2i?

OpenStudy (dtan5457):

@ganeshie8

OpenStudy (jhannybean):

no, it's not complex.

OpenStudy (jhannybean):

This is what you do: \[\frac{5}{3-\sqrt{2}}\cdot \frac{3+\sqrt{2}}{3+\sqrt{2}}\]\[= \frac{5(3+\sqrt{2}}{3^2 - (\sqrt{2})^2}\]NOw simplify this.

OpenStudy (jhannybean):

Meant to write \(5(3+\sqrt{2})\) on top.